A fan motor uses 6630 J of energy in 8.80 s. This energy is used to accelerate a 12.9 kg blade to 26.5 m

s


ms in 7.75 s. What is the efficiency of the fan? Leave your answer as a decimal.


1
Expert's answer
2021-03-08T08:23:46-0500

work done in accelerating the fan is equal to change in it's "K.E=1/2\times 12.9\times26.5^2"

"W=4529.5125J"

rate of work done "=W\/t=4529.5125\/7.75=584.45w"

rate of work done by motor"=P_{in\n}= 6630\/8.8=753.41w"

"efficiency= \frac{P_{out}}{P_{in}}\times100 \%=\frac{584.45}{753.41}\times 100\%=77.6\%"